OpenStudy (anonymous):

If 37.4 grams of water decomposes at 297 Kelvin and 1.30 atmospheres, how many liters of oxygen gas can be produced? Show all of the work used to solve this problem. 2 H2O (l) yields 2 H2 (g) + O2 (g)

OpenStudy (somy):

use PV=nRT

OpenStudy (somy):

\[PV = \frac{ mass }{ Mr }\times RT\]

rishavraj (rishavraj):

18.69

OpenStudy (somy):

\[PV \times Mr = mass \times RT\] make volume as a subject \[V =\frac{ mass \times RT }{ P \times Mr }\]

OpenStudy (somy):

Mr value of O2 gas u can find from periodic table. Covert unit of pressure to Pascals then substitute all the values to the final formula your answer will be in m3 so convert it to dm3 (which is Liters) by multiplying it to 1000

OpenStudy (somy):

@rishavraj please do not give answers. It's against OS rules.

OpenStudy (anonymous):

Would I be using the mole ratio?

OpenStudy (somy):

no you don't need to, since they gave you mass)

rishavraj (rishavraj):

okay

OpenStudy (somy):

by the way u r given Mass of water, so do make sure to find mass of O2 first

OpenStudy (somy):

and yes in that that case use mole ratio @yesy7

OpenStudy (anonymous):

I'm really confused. Okay so I use PV = nRT?

OpenStudy (somy):

sorry i didn't read that the mass was for water forget what i said earlier :) i'll take u step by step first start with finding mole of water

OpenStudy (anonymous):

Okay. Can you help me with that?

OpenStudy (somy):

2 H2O (l) yields 2 H2 (g) + O2 (g) you are given a mass of water so find mole of it \[mole =\frac{ mass }{ Mr }\]

OpenStudy (somy):

u can find Mr right?

OpenStudy (anonymous):

I know R but yeah I'm confused.

OpenStudy (somy):

Mr is molecular mass forget about R for now :D

OpenStudy (somy):

Mr of H2O is = 2 H atoms= 2 1 O atom= 16

OpenStudy (somy):

16+2=18 and since coefficient of H2O in the reaction is 2 then multiply 18 by 2 and this will be your total Mr

OpenStudy (somy):

do u get it so far?

OpenStudy (anonymous):

Oh, I remember that now.

OpenStudy (anonymous):

So Mr is 36?

OpenStudy (somy):

:) good yeah)

OpenStudy (somy):

so now find the mole

OpenStudy (anonymous):

To find the mole I'm dividing 37.4 by 36, right?

OpenStudy (somy):

yes)

OpenStudy (anonymous):

So it would 1 or 1.04

OpenStudy (somy):

1.04 is right ) don't round of too much

OpenStudy (somy):

okay so now if you look at ratio between H2O and O2 what can u say?

OpenStudy (anonymous):

is it 2:1

OpenStudy (somy):

good so what will be O2's mole then?

OpenStudy (somy):

if 2- 1.04 1- ?

OpenStudy (anonymous):

.5?

OpenStudy (somy):

always take 3 significant figures in your answer to be on the safest side

OpenStudy (somy):

0.52 mole

OpenStudy (anonymous):

Okay.

OpenStudy (somy):

so now we have mole of oxygen gas

OpenStudy (somy):

now we can use PV=nRT

OpenStudy (somy):

P is ur pressure V is volume n is mole R is constant = 8.31 T is temperature

OpenStudy (anonymous):

Which mole would I use?

OpenStudy (somy):

first convert pressure to pascals

OpenStudy (somy):

the question is asking u for volume of oxygen so what do u think which mole u have to use?

OpenStudy (anonymous):

So would it be PV = (0.54)(0.0821)(297)

OpenStudy (somy):

mole is 0.52 use 8.31 as your constant

OpenStudy (somy):

because i'm not really familiar with this one, so we can make mistake with units

OpenStudy (anonymous):

sorry typo okay so ' PV = (0.52)(8.31)(297)? Do I just multiply

OpenStudy (somy):

yes

OpenStudy (somy):

afterwards divide the value u get by pressure do make sure to convert atm to Pa

OpenStudy (somy):

1 atm is 101 325 pascals 1.30 atm is ?

OpenStudy (anonymous):

1283.3 / 131.69 = 0.97

OpenStudy (anonymous):

?

OpenStudy (somy):

1283.4/ 131722.5 =

OpenStudy (anonymous):

0.0097

OpenStudy (somy):

then straightaway multiply the answer by 1000 so that u get the answer in dm3

OpenStudy (somy):

when u do it in calculator, do the division, and do not erase it and straightaway multiply by 1000

OpenStudy (anonymous):

so it's 0.9743

OpenStudy (somy):

im getting 9.74 after rounding of

OpenStudy (somy):

in these type of calculations i would suggest you to round of ONLY final answer, so do not round of while you are doing same calculation (meaning u are working in same formula)

OpenStudy (anonymous):

I got the same thing the first time when I did 1283.3/131.69

OpenStudy (somy):

oh that's because u took pascals like that :)

OpenStudy (anonymous):

Oh, that's what I had in my notes so yeah.

OpenStudy (somy):

it's very important to consider units in all calculations thats why i use Pa and u used kPa :)

OpenStudy (anonymous):

I understand.

OpenStudy (somy):

but again if you used kPa make sure your unit of final product matches :)

OpenStudy (somy):

so do u get everything?

OpenStudy (anonymous):

Yeah, it's much clearer now. Thank you for your help!

OpenStudy (somy):

you are welcome :) if you have any doubts feel free to ask :) since i messed up ur brain in the beginning by reading the question wrong lol

OpenStudy (anonymous):

It's fine. I'm doing some more problems so I might have you double check them for me later. :)

OpenStudy (somy):

no problem :)